DX LUCI2 AO Commissioning

02:20
  • Colimated
  • RA: 7:03:45.0327052937212
  • DEC: -12:26:50.7167010355326
02:42
  • New Preset, same target. WFS was not properly sending frames, it was restarted.
  • Bin 1
  • 990 Hz
  • 400 Modes
  • Failed in Optimize Gain
02:48
  • New Preset
  • Bin 1
  • 990Hz
  • 400 Modes
  • Failed in Optimize Gain again
03:13
  • New-New Preset
  • Bin 1
  • 990 Hz
  • 400 Modes
  • Gain 1.34 0.70 0.61
  • Magnitude 8.3
  • Data: 20170321_031629
  • GOPT gain 0.20
03:28
  • Moved Focus (bayside Z) from 55.0 to 55.7, As measured last spring.
  • No obvious improvement, moved back to 55.0
  • Bumped gain down to 0.25 0.10 0.05 (After GOPT)
  • Dataset: 20130321_033311
  • Gain to 0.25 0.10 0.10
  • Dataset: 20130321_033613
  • Magnitude varying wildly due to clouds, now 8.8
03:43
  • New preset to comparison target
  • RA: 7:08:1.2615781591667
  • DEC: -12:24:9.62483581843244
  • Gain after GOPT 0.25 0.10 0.10
  • Dataset: 20170321_034959
  • Bin 1
  • 990Hz
  • 400 Mode
  • Magnitude 7.6
  • Correction was exceedingly poor.
  • Luci2 image 0023
04:08
  • Seeing about 1.26
  • New preset for a quick test
  • AO373
  • RA: 10:02:8.22775385983007
  • DEC: +33:02:20.3539349053545
  • No GOPT
  • LUCI Frame 34
  • Dataset 20170321_042519
  • Gain 1.19 0.91 0.82
  • New Gain 1.15 0.90 0.80
  • GOPT On
  • GOPT ratio: 0.39
  • LUCI frame 035
  • Dataset: 20170321_043045
  • NCPA On
  • LUCI frame 36
  • Dataset: 20170321_043543
  • Open Loop
  • LUCI frame 037
04:30
  • New Flat
  • 20170321_400_1.sav saved as default
  • Old flat pointer: LUCIFER.sav -> 20170304_200900_flao400modes.sav
  • Now LUCIFER points to default.sav

-- %USERSIG{GregoryTaylor - 2017-03-21}%
Topic revision: r2 - 21 Mar 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback